Continuous Improvement is a philosophy that seeks to make never-ending improvements to the process of converting inputs into outputs. The Japanese manufacturer, as well as service providers, have long used this concept. Kaizen is the Japanese word for continuous improvement. Whilst there are several definitions of quality, ultimately it is about satisfying needs and expectations. TQM attempts to maximize the organization’s performance through continual improvement of its products, services, processes or people. TQM supports an organization’s drive to become globally competitive. The future of quality will have a significant impact on global competition, managing customer expectations and dealing with limited resources.

Area of Operations Choose an item of inventory that is used by your organization.
Inventory Management 3.1 Calculate the Economic Order Quality for the selected inventory item.
3.2 Determine the number of times the item needs to ordered.
3.3 Calculate the length of the order cycle.
3.4 Work out the re-order point.
Answer #1

Inventory Management

Welding rods are one of the most important consumables in any mechanical industry. TSAL uses on an average of 10000 units of arc welding rods annually. The ordering cost for one order is $ 2000. This includes the administrative costs and transportation as the supplier is near to the company. The per unit cost is $20 and the holding cost of 20% annually.

The optimal order quantity can be found out by the Economic order quantity formula

3.1 EOQ= Sqrt(2*10000*2000/(0.2*20))

= 3162 units

3.2 Number of times order has to be made = 10000/3162= 3.16 times

3.3 Length of order cycle considering 365 days in a year

=365/3.16= 115.5 days

3.4 Reorder point

The lead time is 2 days

ROP = 10000*2/365= 54.79= 55 units

When the inventory level reaches 55 units make an order of 3162 units

Please comment if any doubt

Rate me

Thanks